A rectangle plot of land is to be fenced off so that the area enclosed will be 400 ft2

. Let 

L

be the 

length of fencing needed and 

x

the length of one side of the rectangle. Show that

L ="2x" + (800/x)

for 𝑥 > 0. What is the shortest possible length of such a fence?


1
Expert's answer
2021-02-24T06:48:56-0500

Let x be one side of a rectangle and y the other. Then the area of a rectangle is, obviously, "S=xy=400\text{ ft}^2" and thus "y = \frac{400}{x}", x and y are both ">0". Now the perimeter of this rectangle (and thus the total length of fence needed) is given by "L=2x+2y = 2x + 2\cdot \frac{400}{x}=2x+\frac{800}{x}".

Now to find the minimum of a total length we will calculate the derivative of L with respect to x : "\frac{dL}{dx} =2-\frac{800}{x^2}" and the zero of this derivative is "x_0 = 20". Thus the minimal length is "L = 2\cdot 20+\frac{800}{20} = 80 \text{ ft}".